OpenStudy (anonymous):

Plz i need urgent help~~Giving out a medal and fanning anyone who helps thank you!~~ Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x2 + 13x + 15 2x2 + 27x + 13 5x2 − 2x − 7 10x2 − 7x − 3 8x2 − 11x + 3

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

Hey carroll do you think you could help me?

OpenStudy (anonymous):

Let's do 8x2 − 11x + 3 First what you do is multiple 8 and 3 together . Which should look like this \[8\times3\] Which equals to 24! After you have done that you find TWO numbers . You can multiply to 24 and you add -11 to it and your answer will be\[-8\times-3\] The we have to rewrite the the equation as follow \[8\times2-8\times-3\times +3\] You factor the first two terms together and the second two terms together \[8\times(\times-1)-3(\times-1)\] The factor this by taking \[(\times-1)\] out as follow \[(\times-1)(8\times-3)\] Understand?

OpenStudy (anonymous):

Ok i'm kinda getting it but i'm still a little confused

OpenStudy (anonymous):

where r u confused?

OpenStudy (anonymous):

Im not sure if shes giving me the factored answer and i'm not sure what to do with the next question because i dont know what is the steps

OpenStudy (anonymous):

I pretty much kinda did your work for you sweet heart.

OpenStudy (anonymous):

no problem @HelloGoodmorning if problem persist come back to us.

OpenStudy (anonymous):

Oh but could you explain how to solve one you dont have to solve mines please just give me an example because i have to do to of them unfortunately

OpenStudy (anonymous):

Thanks heybhai

OpenStudy (anonymous):

Unfortunately i still need help

OpenStudy (precal):

ok I have another way to do it

OpenStudy (precal):

it is called the slide and divide method so let us slide 2 to the back by multiplying it to 15

OpenStudy (precal):